solve the system of equations by the addtion method 3x-4y=7
5x+8y=19

Answer by Alan3354(21608)

3x-4y=7
5x+8y=19
6x-8y=14 1st eqn x 2
-------- Add
11x = 33
x = 3
-------
y = 1/2